A single-lane street, 12 feet wide goes through a semi-circular tunnel with a radius of 10 feet. How high is the tunnel at the edge of each lane?


1
Expert's answer
2021-11-16T13:54:57-0500

Let "r=10" ft, "l=12" ft.

"h=\sqrt{r^2-(\frac l2)^2}=\sqrt{10^2-6^2}=8" ft.